This is Python Coding question.

Can anyone help me figuring these out?

(Please don't use built-in function, and leave comments for my understanding)

Thanks!!

1. Write a function called span(n) that takes as a parameter a 2D list of numbers called n and returns the difference between the largest and smallest number in n. For example, it should return the value 6 for the list [[2, 1, 3],[3, 7, 4]] since the largest number is 7 and the smallest is 1, and their difference is 6.

--> (topic is 2D list)

2. A mortgage loan is charged interest every month based on the current balance of the loan. If the annual rate is r%, then the interest for the month is r/12% of the current balance. After the interest is applied each month, the borrower makes a payment which reduces the amount owed. Write a function mortgage(principal, rate, years, payment) that prints out the balance at the end of each month for years number of years.

3. Write a function that takes an integer parameter n and returns the nth Fibonacci number. The Fibonacci numbers start 1,1,2,3,5,8,... and every subsequent one is the sum of the previous two.
Answer #1

Output

Code with comments

def span(n): #span function
mini=1000000 #minimum number in list
maxi=-100000 #maxiim number in list
for i in n: #for each list i in 2d list n
for j in i: #for each element in i
mini=min(j,mini) #update minimum
maxi=max(j,maxi) #update maximum
print("Minimum:",mini)
print("Maximum:",maxi)
print("Span output: "+str(maxi-mini))

span( [[2, 1, 3],[3, 7, 4]] ) #test

def mortgage(principal,rate,year,payment): #mortgage function
print("\nMortgage balance")
month=0 #month count
print("Month: "+str(month)+", Principal: "+str(principal)) #initial
while principal>0 and month<12*year: #either loan is paid or number of years
interest=principal*(rate/12) #calculate interest
amount=payment-interest #principal paid each month
principal=principal-amount #deduct from total
month=month+1
print("Month: "+str(month)+", Principal: "+str(principal))

mortgage(1000,0.12,2,40) #test

def nthfib(n): #nth fibonacci
if n<=2: # 1st and 2nd are 1
return 1
a=1
b=1
for i in range(3,n+1): #starting from 3rd to nth
c=a+b
a=b
b=c
return c
print("\n7th fibonacci number is: "+str(nthfib(7))) #test

For indentation purpose